fundamental frequency of pipe

A certain organ pipe, open at both ends, produces a fundamental frequency of 271 Hz in air.
If the pipe is filled with helium at the same temperature, what fundamental frequency f He will it produce? Take the molar mass of air to be 28.8 g/mol and the molarmass of helium to be 4.00g/mol .
Answer #1

f is proportional to velocity

V = [RT/M]^0.5

Sof1/f2 =√M2/√M1 =√4/√28.8 =0.37267

So, f2 = f1/0.37267 =727.169 Hz

answered by: LaScott Ellis